OpenStudy (wolfadaskies):

I just need a bit of help with these few Alg. 1 questions, please

OpenStudy (anonymous):

what are the questions?

OpenStudy (wolfadaskies):

SImplify the radical expression: (sq. root)100h^10

OpenStudy (wolfadaskies):

Simplify the rational expression: (sq. root)8h^10

OpenStudy (anonymous):

like this? \[\sqrt{100*h^{10}}\]

OpenStudy (wolfadaskies):

Yeah! Sorry, I don't know the commands for format here yet lol

OpenStudy (anonymous):

ok, so for the first one, you can pull out the square root of 100. do you know what that is?

OpenStudy (wolfadaskies):

10

OpenStudy (anonymous):

correct, now you can pull out the square root of h^10 (divide the exponent by 2)

OpenStudy (wolfadaskies):

3, but what do I do with the exponent?

OpenStudy (anonymous):

actually it would be h^5 when you pull it outside the square root

OpenStudy (wolfadaskies):

Or 5, my bad

OpenStudy (wolfadaskies):

so 10h^5 ?

OpenStudy (anonymous):

Yep! thats exactly right! good job! :)

OpenStudy (anonymous):

the second one can be done the same way

OpenStudy (wolfadaskies):

Ohh, I see, thank you! Could you help me with a few more?

OpenStudy (whpalmer4):

Remember, when you multiply two exponentials with the same base, you get the new exponent by adding them together. \[x^a*x^b = x^{a+b}\] So if your answer was \(10h^5\), squaring it should give you what you started with under the radical sign. \[10h^5*10h^5 = 10*10*h^5*h^5=100*h^5*h^5=100*h^{5+5} = 100*h^{10}\checkmark\]

OpenStudy (whpalmer4):

Gahh. stupid OpenStudy window sizing... \[=100*h^{5+5} = 100*h^{10}\]

OpenStudy (anonymous):

@whpalmer4 he/she is actually trying to find the square root of (100*h^10)

OpenStudy (wolfadaskies):

Ah, I understand a bit better now. So with this: V45 would it be 5V3 or 3V5 ?

OpenStudy (anonymous):

yes @Wolfadaskies thats right, good job!

OpenStudy (wolfadaskies):

(thats supposed to be a sq. root sign, sorry)

OpenStudy (wolfadaskies):

Is either on e of those formats correct, then?

OpenStudy (anonymous):

i assumed you meant square root but couldnt figure out the formatting but it is supposed to be a square root so \[3\sqrt{5}\]

OpenStudy (wolfadaskies):

AH yeah thats what I meant to put ^_^/ thank you

OpenStudy (anonymous):

sure thing! :) keep up the good work!

OpenStudy (wolfadaskies):

Thank you so much :)